Show that | z + w |^2 − | z − w |^2 = 4Re (zŵ).


1
Expert's answer
2021-12-27T18:19:56-0500

Let "z=a+bi, w=c+di, a,b,c,d\in\R."

Then


"|z+w|^2=(a+c)^2+(b+d)^2"

"|z-w|^2=(a-c)^2+(b-d)^2"

"|z+w|^2-|z-w|^2=(a+c)^2+(b+d)^2"

"-((a-c)^2+(b-d)^2)=4ac+4bd"

"=4(ac+bd)"

"z\u0175=(a+bi)(c-di)=ac+bd+(-ad+bc)i"

"Re(z\u0175)=ac+bd"

Hence


"|z+w|^2-|z-w|^2=4(ac+bd)=4Re(z\u0175),"

"a,b,c,d\in \R"

Therefore


"|z+w|^2-|z-w|^2=4Re(z\u0175)"